A local brokerage firm is offering a zerominuscoupon certificate of deposit for $10,000. At maturity, three years from now, the investor will receive $14,000. What is the rate of return on this investment?
Answer #1

Ans 11.87%

FV = Future Value
PV = Present Value
r = rate of interest
n= no of period
FV/ PV = (1 + r )^n
14000 / 10000 = (1 + r%)^3
1.4 = (1 + r%)^3
r = 11.87%